what is the answer to log (x+2) - logx = 3?
Answer by user_dude2008(1862)

log (x+2) - log(x) = 3


log [(x+2)/x] = 3


(x+2)/x = 10^3


(x+2)/x = 1000


x+2 = 1000x


999x=2


Ans: x = 2/999